# Towards Understanding the Working Mechanism of Text-to-Image Diffusion Model

### v1

@@ 1 Introduction
In real-world application, the Text-to-Image (T2I) generation has long been explored owing to its wide applications [46; 31; 32; 2; 15], whereas the Diffusion Probabilistic Model (DPM) [12; 37; 32] stands out as a promising approach, thanks to its impressive image synthesis capability. Technically, the DPM is a hierarchical denoising model, which gradually purifies noisy data from a standard Gaussian to generate an image. In the existing literature [31; 34; 28; 6], the framework of (latent) Stable Diffusion model [31] is a backbone technique in T2I generation via DPM. In this approach, the text prompt is encoded by a CLIP text encoder [30], and injected into the diffusion image decoder as a condition to generate a target image (latent encoded by VQ-GAN in [31]) that is consistent with the text prompt. Though this framework works well in practice, the working mechanism behind it, especially for the text prompt, remains to be explored. Therefore, in this paper, we systematically explore the working mechanism of stable diffusion.
@@
Following the phenomenons, we investigate the effect of encoded tokens in the text prompt of T2I generation during the two stages, where each token is encoded by an auto-regressive CLIP text encoder. The text prompt has a length of 76, and is enclosed by special tokens [SOS] and [EOS], 111[EOS] contains the overall information in text prompt due to the auto-regressive CLIP text-encoder at the beginning and end of the text prompt, respectively. Therefore, we categorize the tokens into three classes, i.e., [SOS], semantic tokens, and [EOS]. Notably, the special token [SOS] does not contain information, due to the auto-regressive encoding of the text prompt. Thus, our investigations into the influence of tokens will primarily focus on the semantic tokens and [EOS]. Surprisingly, we find that compared with semantic tokens, the special token [EOS] has a larger impact during generation.
@@ 2 Related Work
In this paper, our exploration is based on the Stable Diffusion [32], which now terms to be a standard T2I generation technique based on DPM [12; 36], and has been applied into various computer vision domains e.g., 3D [29; 33; 19] and video generation [26; 4]. In practice, the goal of T2I is generating an image that is consistent with a given text injected into the cross-attention module [42] of the image decoder. Therefore, understanding the working mechanism of stable diffusion potentially improves the existing techniques [1]. Unfortunately, to the best of our knowledge, the problem is limited explored, expected in [45; 35], where they similarly observe the low-frequency signals are firstly recovered in the denoising process. However, further explanations for this phenomenon are neglected in these works.
@@ 3 Preliminaries
Finally, the diffusion model (usually UNet) takes the text prompt as input to the cross-attention module in each basic block  [^2] The model is stacked by such basic blocks sequentially with residual module, self-attention module, and cross-attention module in each of it. of diffusion model with output ${\rm Attention}(Q,K,V)=\mathrm{Softmax}(QK^{\top}/\sqrt{d})V$ ($d$ is dimension of image feature), where $\phi(\boldsymbol{x}_{t})$ is the feature of image, and
@@ 4.2 Frequency Analysis
For all $u\in[M],v\in[N]$, with probability at least $1-\delta$, we have
$\small\left\|F_{\boldsymbol{\epsilon}_{t}}(u,v)\right\|^{2}\leq\frac{1}{\sqrt{MN}}\left(1+\sqrt{8\log{\frac{2MN}{\delta}}}\right).$ (7)
@@
However, as visualized in Figure LABEL:fig:corrupted_data  [^3] Please note that the $t$ in Figure LABEL:fig:corrupted_data and the other parts of this paper refers to the corresponding $t$-th time step of 50-steps DDIM sampling, which corresponds to $20t$ steps in [12], in contrast to high-frequency part of image, the image’s low-frequency parts  [^4] We distinguish low frequency by threshold 20%, i.e., the lowest 20% parts of spectrum are low-frequency. are more robust than the ones of high-frequency. For example, for $t=20$ in Figure LABEL:fig:corrupted_data, the shape of the clock is still perceptible in the low-frequency part of the image.  [^5] This fact is also observed in [35; 45], while they leverage this phenomenon to refine the architecture of UNet, and the rationale behind the phenomenon is unexplored. If this fact is generalized to image latent, then it explains the two stages of generation as observed in Section 4.1. Because the low-frequency parts are not corrupted until the end of the adding noise process. Then, it will be recovered at the beginning of the reverse denoising process.
@@ 5 The Working Mechanism of Text Prompt
In T2I diffusion, the text prompt is encoded by auto-regressive CLIP text encoder, with semantic tokens (SEM) enclosed with special tokes [SOS] and [EOS]. For such three classes of tokens, as the information in these tokens is conveyed by the cross-attention module, we first compute the averaged weights over pixels in the cross-attention map for each class. The weights are computed by taking the average over PromptSet and presented in Figure 3. As can be seen, the weights of [SOS] are significantly larger than the other classes. However, due to the CLIP text encoder is an auto-regressive model, [SOS] does not contain any semantic information. Therefore, we conclude that the influence of [SOS] is mainly adjusting the whole cross-attention map i.e., weights on the other tokens. A similar phenomenon is observed in single-modality LLM [44]. As the information of text prompt is conveyed by semantic tokens and [EOS], we will focus on them instead of [SOS] in the sequel.
@@
We examine the generated images under these artificially constructed text prompts (namely Switched-PromptSet (S-PromptSet)). We call $A$ from PromptA as “source” and $B$ from [EOS]B as “target” for “[SOS] + Prompt $A$ + [EOS]B”, and vice versa. For the generated images under these prompts, we measure their alignments with the source and target prompts, respectively. The used metrics are the three standard ones in measuring text-image alignment: CLIPScore [30; 10], BLIP-VQA [18; 14], and MiniGPT4-CoT [49; 14] (details are in Appendix B).
@@ Discussion.
Comparing (3) and (8), it holds $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)\propto\nabla_{\boldsymbol{x}}\log{p_{t}(\boldsymbol{x}_{t})}$  [^6] This can be verified by the training strategy of it in [31], where $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)$ is used to predict noise in noisy data without condition injected. and $w(\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C})-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset))\propto\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$. From [37], the denoising process (4) aims to maximize log-likelihood $\log{p_{0}(\boldsymbol{x}_{0}\mid\mathcal{C})}$. Then, moving along the direction $\nabla_{\boldsymbol{x}}\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$ (leads to large $\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$) push $\boldsymbol{x}_{t}$ to be aligned with the text prompt $\mathcal{C}$ during the decreasing of $t$. Adding such a moving direction is standard in conditional generation [25; 36; 8; 24]. As shown in Figure LABEL:fig:noise_norm, during the denoising process, $\boldsymbol{x}_{t}$ will gradually to be consistent with $\mathcal{C}$, so that $\nabla\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$ will decrease with $t$. Thereafter, we observe the impact of text prompt conveyed by this term decreases with $t\to 0$. Notably, owing to the quickly reconstructed overall shape of image in Section 4, the generated $\boldsymbol{x}_{t}$ will quickly be consistent with $\mathcal{C}$, so that explain the quickly decreasing $\nabla\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$.
@@ Appendix A Proofs of Proposition 1
with $\theta^{kl}_{uv}$ is the $kl$-th angle in complex value space, and we may simplify it as $\theta^{kl}$ for ease of notations.
@@
$\left\|F_{\boldsymbol{\epsilon}}(u)\right\|^{2}=F_{\boldsymbol{\epsilon}}(u)\overline{F_{\boldsymbol{\epsilon}}(u)}=\frac{1}{M^{2}}\boldsymbol{\epsilon}^{\top}\Lambda\textbf{1}\textbf{1}^{\top}\bar{\Lambda}\boldsymbol{\epsilon},$ (11)
where $\Lambda=\mathsf{diag}(e^{-i\theta^{0}},\cdots,e^{-i\theta^{M-1}})$. Then let $\boldsymbol{P}=(\sqrt{1/M}\textbf{1}^{\top},\cdots,)^{\top}\bar{\Lambda}$, where $(\sqrt{1/M}\textbf{1}^{\top},\cdots,)^{\top}$ is constructed by vector $\sqrt{1/M}\textbf{1}$ and its orthogonal complement. We can verify that $\boldsymbol{P}$ is an orthogonal matrix. Then, let $\boldsymbol{y}=\boldsymbol{P}\boldsymbol{\epsilon}$, so that $\boldsymbol{y}$ has the same distribution with $\boldsymbol{\epsilon}$. Thus
@@
which proves our conclusion. ∎
@@ Appendix D Conveyed Information in Semantic Tokens
Appendix E More Evidences on [EOS] Contains More Information
@@
Appendix F Key or Value Dominates the Influence?
@@
Appendix G More Generated Images
In this section, we present more generated images.
G.1 [EOS] Substitution
@@ Appendix G More Generated Images
G.2 Generated Images in Paragraph “Acceleration of Sampling” of Section 6
In this subsection, we present more generated images under noise prediction (9) with varied $a$. The results are in Figure 17 and 18.

### v2

@@ Abstract
**footnotetext: equal contributionfootnotetext: corresponding to Mingyang Yi: yimingyang@ruc.edu.cn
@@ 1 Introduction
In real-world application, the Text-to-Image (T2I) generation has long been explored owing to its wide applications [47; 31; 32; 2; 15], whereas the Diffusion Probabilistic Model (DPM) [12; 37; 32] stands out as a promising approach, thanks to its impressive image synthesis capability. Technically, the DPM is a hierarchical denoising model, which gradually purifies noisy data from a standard Gaussian to generate an image. In the existing literature [31; 34; 28; 6], the framework of (latent) Stable Diffusion model [31] is a backbone technique in T2I generation via DPM. In this approach, the text prompt is encoded by a CLIP text encoder [30], and injected into the diffusion image decoder as a condition to generate a target image (latent encoded by VQ-GAN in [31]) that is consistent with the text prompt. Though this framework works well in practice, the working mechanism behind it, especially for the text prompt, remains to be explored. Therefore, in this paper, we systematically explore the working mechanism of stable diffusion.
@@
Following the phenomenons, we investigate the effect of encoded tokens in the text prompt of T2I generation during the two stages, where each token is encoded by an auto-regressive CLIP text encoder. The text prompt has a length of 76, and is enclosed by special tokens [SOS] and [EOS], [EOS] contains the overall information in text prompt due to the auto-regressive CLIP text-encoder at the beginning and end of the text prompt, respectively. Therefore, we categorize the tokens into three classes, i.e., [SOS], semantic tokens, and [EOS]. Notably, the special token [SOS] does not contain information, due to the auto-regressive encoding of the text prompt. Thus, our investigations into the influence of tokens will primarily focus on the semantic tokens and [EOS]. Surprisingly, we find that compared with semantic tokens, the special token [EOS] has a larger impact during generation.
@@ 2 Related Work
In this paper, our exploration is based on the Stable Diffusion [32], which now terms to be a standard T2I generation technique based on DPM [12; 36], and has been applied into various computer vision domains e.g., 3D [29; 33; 19] and video generation [26; 4]. In practice, the goal of T2I is generating an image that is consistent with a given text injected into the cross-attention module [42] of the image decoder. Therefore, understanding the working mechanism of stable diffusion potentially improves the existing techniques [1]. Unfortunately, to the best of our knowledge, the problem is limited explored, expected in [46; 35], where they similarly observe the low-frequency signals are firstly recovered in the denoising process. However, further explanations for this phenomenon are neglected in these works.
@@ 3 Preliminaries
Finally, the diffusion model (usually UNet) takes the text prompt as input to the cross-attention module in each basic block The model is stacked by such basic blocks sequentially with residual module, self-attention module, and cross-attention module in each of it. of diffusion model with output ${\rm Attention}(Q,K,V)=\mathrm{Softmax}(QK^{\top}/\sqrt{d})V$ ($d$ is dimension of image feature), where $\phi(\boldsymbol{x}_{t})$ is the feature of image, and
@@ 4.2 Frequency Analysis
For all $u\in[M],v\in[N]$, with high probability, the complex number $F_{\boldsymbol{\epsilon}_{t}}(u,v)$ satisfies
$\small\left\|F_{\boldsymbol{\epsilon}_{t}}(u,v)\right\|^{2}\approx\mathcal{O}\left(\frac{1}{MN}\right).$ (7)
@@
However, as visualized in Figure LABEL:fig:corrupted_data Please note that the $t$ in Figure LABEL:fig:corrupted_data and the other parts of this paper refers to the corresponding $t$-th time step of 50-steps DDIM sampling, which corresponds to $20t$ steps in [12], in contrast to high-frequency part of image, the image’s low-frequency parts We distinguish low frequency by threshold 20%, i.e., the lowest 20% parts of spectrum are low-frequency. are more robust than the ones of high-frequency. For example, for $t=20$ in Figure LABEL:fig:corrupted_data, the shape of the clock is still perceptible in the low-frequency part of the image. This fact is also observed in [35; 46], while they leverage this phenomenon to refine the architecture of UNet, and the rationale behind the phenomenon is unexplored. If this fact is generalized to image latent, then it explains the two stages of generation as observed in Section 4.1. Because the low-frequency parts are not corrupted until the end of the adding noise process. Then, it will be recovered at the beginning of the reverse denoising process.
@@ 5 The Working Mechanism of Text Prompt
In T2I diffusion, the text prompt is encoded by auto-regressive CLIP text encoder, with semantic tokens (SEM) enclosed with special tokes [SOS] and [EOS]. For such three classes of tokens, as the information in these tokens is conveyed by the cross-attention module, we first compute the averaged weights over pixels in the cross-attention map for each class. The weights are computed by taking the average over PromptSet and presented in Figure 3. As can be seen, the weights of [SOS] are significantly larger than the other classes. However, due to the CLIP text encoder is an auto-regressive model, [SOS] does not contain any semantic information. Therefore, we conclude that the influence of [SOS] is mainly adjusting the whole cross-attention map i.e., weights on the other tokens. To further verify this conclusion, we conduct experiments in Appendix I.1. A similar phenomenon is observed in single-modality LLM [45]. As the information of text prompt is conveyed by semantic tokens and [EOS], we will focus on them instead of [SOS] in the sequel.
@@
We examine the generated images under these artificially constructed text prompts (namely Switched-PromptSet (S-PromptSet)). We call $A$ from PromptA as “source” and $B$ from [EOS]B as “target” for “[SOS] + Prompt $A$ + [EOS]B”, and vice versa. For the generated images under these prompts, we measure their alignments with the source and target prompts, respectively. The used metrics are the three standard ones in measuring text-image alignment: CLIPScore [30; 10], BLIP-VQA [18; 14], and MiniGPT4-CoT [51; 14] (details are in Appendix B).
@@ Discussion.
Comparing (3) and (8), it holds $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)\propto\nabla_{\boldsymbol{x}}\log{p_{t}(\boldsymbol{x}_{t})}$ This can be verified by the training strategy of it in [31], where $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset)$ is used to predict noise in noisy data without condition injected. and $w(\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\mathcal{C})x-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(t,\boldsymbol{x}_{t},\emptyset))\propto\nabla\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$. From [37], the denoising process (4) aims to maximize log-likelihood $\log{p_{0}(\boldsymbol{x}_{0}\mid\mathcal{C})}$. Then, moving along the direction $\nabla_{\boldsymbol{x}}\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$ (leads to large $\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$) push $\boldsymbol{x}_{t}$ to be aligned with the text prompt $\mathcal{C}$ during the decreasing of $t$. Adding such a moving direction is standard in conditional generation [25; 36; 8; 24]. As shown in Figure LABEL:fig:noise_norm, during the denoising process, $\boldsymbol{x}_{t}$ will gradually to be consistent with $\mathcal{C}$, so that $\nabla\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$ will decrease with $t$. Thereafter, we observe the impact of text prompt conveyed by this term decreases with $t\to 0$. Notably, owing to the quickly reconstructed overall shape of image in Section 4, the generated $\boldsymbol{x}_{t}$ will quickly be consistent with $\mathcal{C}$, so that explain the quickly decreasing $\nabla\log{p_{t}(\mathcal{C}\mid\boldsymbol{x}_{t})}$.
Remark 2.
In this section, we verify the injected textual information are all conveyed in the first stage of diffusion process. In fact, this phenomenon is also generalized to the other types of information, e.g., conditional image information in subject-driven generation [44; 50], we verify this in Appendix 11
@@ Appendix A Proofs of Proposition 1
with $\theta^{kl}_{uv}$ is the $(k,l)$-th angle in complex value space, and we may simplify it as $\theta^{kl}$ for ease of notations.
@@
$\left\|F_{\boldsymbol{\epsilon}}(u)\right\|^{2}=F_{\boldsymbol{\epsilon}}(u)\overline{F_{\boldsymbol{\epsilon}}(u)}=\frac{1}{M^{2}}\boldsymbol{\epsilon}^{\top}\Lambda\textbf{1}\textbf{1}^{\top}\bar{\Lambda}\bar{\boldsymbol{\epsilon}},$ (11)
where $\Lambda=\mathsf{diag}(e^{-i\theta^{0}},\cdots,e^{-i\theta^{M-1}})$. Then let $\boldsymbol{P}=(\sqrt{1/M}\textbf{1}^{\top},\cdots,)^{\top}\bar{\Lambda}$, where $(\sqrt{1/M}\textbf{1}^{\top},\cdots,)^{\top}$ is constructed by vector $\sqrt{1/M}\textbf{1}$ and its orthogonal complement. We can verify that $\boldsymbol{P}$ is an orthogonal matrix. Then, let $\boldsymbol{\epsilon}=\boldsymbol{P}^{\top}\boldsymbol{y}$, so that $\boldsymbol{y}$ has the same distribution with $\boldsymbol{\epsilon}$. Thus
@@
which proves that the target $\|F_{\boldsymbol{\epsilon}_{t}}(u,v)\|^{2}$ is of order $\mathcal{O}(\frac{1}{M})$, so that verify our conclusion. ∎
@@ Appendix D Conveyed Information in Semantic Tokens
Figure 13: Generated images with prompts only contain information from [SOS] or [EOS].
@@
Appendix E [SOS] Contains no Textual Information
As mentioned in Section 5.1, the special token [SOS] is supposed to contain no textual information, due to the auto-regressive textual prompt encoder. To further verify this, we conduct the following two types of prompts. 1): all 77 tokens are [SOS] from the given text prompt, 2) except the first [SOS] token, all other 76 tokens are [EOS] from the given text prompt. Then, we generate images under these prompts with SD v1.5. The generated images as in Figure 13. As can be seen, for the first type of prompt, no textual information is conveyed, while the phenomenon disappears in the second type of prompt. This observation verifies our conclusion that [SOS] contains no textual information.
Appendix F More Evidences on [EOS] Contains More Information
@@
Appendix G Key or Value Dominates the Influence?
@@
Figure 17: We implement our sampling strategy on subject-driven generation model, AnyDoor [44]. We remove the condition image from different time steps (denote as $a$) during denoising process. The generated images still preserve the specific details as baseline model (start point $a=0$) when start removing time steps is set to 20.
Figure 18: We implement our sampling strategy on human face generation model, PhotoMaker. We remove the condition (text prompts and reference face) from different time steps (denote as $a$) during denoising process. The generated images and faces still preserve the specific details as baseline model (start point $a=0$) when start removing time steps is set to 20.
@@
Appendix H Firstly Conveyed Information for Subject-Driven Generation
As mentioned in Section 5.1, the injected textual information is conveyed in the first stage of diffusion model. Since the information is conveyed in the cross-attention module of model, we speculate this phenomenon may be generalized to the other conditional generation task, e.g., subject-driven generation with an extra image as condition.
To see this, we conduct experiments under two tasks: zero-shot subject-driven generation and human face generation. For such two tasks, there is extra reference image (given subject and human face) used as condition to guide image generation. We use the sampling strategy as in Figure 10 for the two tasks, respectively follow the backbone methods AnyDoor [44] and Photomaker [50].
As can be seen, the generation results verify the conclusion that for conditional information from in the other modality than text, they will be also firstly conveyed in during the diffusion process.
Appendix I More Generated Images
I.1 [EOS] Substitution
@@ Appendix I More Generated Images
I.2 Generated Images in Paragraph “Acceleration of Sampling” of Section 6
Next, we present more generated images under noise prediction (9) with varied $a$ in Figure 20 and 21.